OpenStudy (anonymous):

Find all real numbers: 3sin^2x-sinx=0 HELP! I just basically need help knowing when i add/sub pi or 2pi

OpenStudy (anonymous):

6sin^2 (x) cos^2 (x) -sinx=0 Take sinx common you get sin(x)=0 As you know sin(n pi) =0 so x= n*pi where n is an integer... For remaining portion 6sinx cos^2 (x)-1=0 we get sinx cos^2 x=1/6 as cos^2 x= 1-sin^2 x sinx -sin^3 x=1/6 solve the cubic equation

OpenStudy (campbell_st):

why not just factor out sin(x) \[\sin(x)[3\sin(x) -1] = 0\] so fins the values where sin(x) = 0 and sin(x) = 1/3 they are the solutions.

OpenStudy (anonymous):

i lknow you take the inverse sin for both but then it tells me to sub tract the inverse of 1/3 from pi....and idk why

OpenStudy (campbell_st):

will if you know the sin curve it looks like  so that covers sin(x) = 0 now for the other angle \[x = \sin^{-1}(\frac{1}{3})\] just type it into a calculator.